# Local certification of geometric graph classes

### v4

@@ Acknowledgement
Note that $\ell-\delta\leq d_{2}(v_{1},v_{\ell+1})\leq\ell$. We place $\ell+1$ evenly spaced points $u_{1},\ldots,u_{\ell+1}$ on the line-segment $[v_{1},v_{\ell+1}]$ with $u_{1}=v_{1}$, $u_{\ell+1}=v_{\ell+1}$. For $12\leq i\leq\ell+$ 1, the distance $d_{2}(u_{i},v_{i})$ is maximized when $i=\left\lceil\frac{\ell+1}{2}\right\rceil$, and is at most the distance between the middle of the line-segment $[v_{1},v_{\ell+1}]$ and a point $c$ at Euclidean distance $\ell/2$ from $v_{1}$ and $v_{\ell+1}$. By Pythagoras’ theorem, this distance is at most $\gamma:=\sqrt{\ell^{2}/4-(\ell-\delta)^{2}/4}=\sqrt{\ell\delta/2}+O(\ell^{-1/2})$, as desired.
For the second part of the lemma, we observe that each vertex $v_{i}$ satisfies $d_{2}(v_{i},v_{1})+d_{2}(v_{i},v_{\ell+1})\leq\ell$, so each such point is contained in the region bounded by the ellipse with foci $v_{1}$ and $v_{\ell+1}$, width $\ell$, and height at most $2\gamma$ (by the preceding paragraph). Let $m$ denote the midpoint of the line segment $[v_{1},v_{\ell+1}]$. Let $p$ be a point on the ellipse, and let $p^{\prime}$ be the projection of $p$ on the line $(v_{1},v_{\ell+1})$. By the standard description of an ellipse in Cartesian coordinates, if $x:=d_{2}(m,p^{\prime})$ then $y:=d_{2}(p,p^{\prime})=\tfrac{2\gamma}{\ell}\sqrt{\ell^{2}/4-x^{2}}$. If we take $(1-\alpha)\ell/2\leq x\leq\ell/2$ for some $\alpha>0$, then $y\leq\gamma\sqrt{1-(1-\alpha)^{2}}=\lambda^{\prime}\gamma$ for $\lambda^{\prime}:=\sqrt{2\alpha-\alpha^{2}}$ independent of $\ell$ and $\delta$. Let $p^{\prime\prime}$ be the point of $[v_{1},v_{\ell+1}]$ at distance $\tfrac{\ell-\delta}{\ell}d_{2}(v_{1},p)$ from $v_{1}$. Then $d_{2}(p^{\prime},p^{\prime\prime})\leq d_{2}(v_{1},p)-\tfrac{\ell-\delta}{\ell}d_{2}(v_{1},p)=\tfrac{\delta}{\ell}d_{2}(v_{1},p)$. It then follows from Pythagoras’ theorem that $d_{2}(p,p^{\prime\prime})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\tfrac{\delta^{2}}{\ell^{2}}d_{2}(v_{1},p)^{2}}$. If $d_{2}(v_{1},p)\leq\alpha\ell$ then $d_{2}(p,p^{\prime\prime})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}\delta^{2}}\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}}$ since $\delta\leq 1$. By considering $p$ above as the location of $v_{i}$ ($1\leq i\leq\alpha\ell$) that maximizes its distance to $u_{i}$ (whose location corresponds to $p^{\prime\prime}$ above), we conclude that $d_{2}(v_{i},u_{i})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}}=\sqrt{(2\alpha-\alpha^{2})\gamma^{2}+\alpha^{2}}$, as desired. ∎
[…]
8

### v5

@@ Acknowledgement
Note that $\ell-\delta\leq d_{2}(v_{1},v_{\ell+1})\leq\ell$. We place $\ell+1$ evenly spaced points $u_{1},\ldots,u_{\ell+1}$ on the line-segment $v_{1},v_{\ell+1}]$ with $u_{1}=v_{1}$, $u_{\ell+1}=v_{\ell+1}$. For $2\leq i\leq\ell+1$, the distance $d_{2}(u_{i},v_{i})$ is maximized when $i=\left\lceil\frac{\ell+1}{2}\right\rceil$, and is at most the distance between the middle of the line-segment $v_{1},v_{\ell+1}]$ and a point $c$ at Euclidean distance $\ell/2$ from $v_{1}$ and $v_{\ell+1}$. See Figure 20 for an illustration. By Pythagoras’ theorem, this distance is at most $\gamma:=\sqrt{\ell^{2}/4-(\ell-\delta)^{2}/4}=\sqrt{\ell\delta/2}+O(\ell^{-1/2})$, as desired.
Figure 20. Illustration of the proof of Lemma 7.9.
For the second part of the lemma, we observe that each vertex $v_{i}$ satisfies $d_{2}(v_{i},v_{1})+d_{2}(v_{i},v_{\ell+1})\leq\ell$, so each such point is contained in the region bounded by the ellipse with foci $v_{1}$ and $v_{\ell+1}$, width $\ell$, and height at most $2\gamma$ (by the preceding paragraph). Let $m$ denote the midpoint of the line segment $v_{1},v_{\ell+1}]$. Let $p$ be a point on the ellipse, and let $p^{\prime}$ be the projection of $p$ on the line $(v_{1},v_{\ell+1})$. By the standard description of an ellipse in Cartesian coordinates, if $x:=d_{2}(m,p^{\prime})$ then $y:=d_{2}(p,p^{\prime})=\tfrac{2\gamma}{\ell}\sqrt{\ell^{2}/4-x^{2}}$. If we take $(1-\alpha)\ell/2\leq x\leq\ell/2$ for some $\alpha>0$, then $y\leq\gamma\sqrt{1-(1-\alpha)^{2}}=\lambda^{\prime}\gamma$ for $\lambda^{\prime}:=\sqrt{2\alpha-\alpha^{2}}$ independent of $\ell$ and $\delta$. Let $p^{\prime\prime}$ be the point of $v_{1},v_{\ell+1}]$ at distance $\tfrac{\ell-\delta}{\ell}d_{2}(v_{1},p)$ from $v_{1}$ (so that if $v_{i}$ coincides with $p$, then $u_{i}$ coincides with $p^{\prime\prime}$). Then $d_{2}(p^{\prime},p^{\prime\prime})\leq d_{2}(v_{1},p)-\tfrac{\ell-\delta}{\ell}d_{2}(v_{1},p)=\tfrac{\delta}{\ell}d_{2}(v_{1},p)$. It then follows from Pythagoras’ theorem that $d_{2}(p,p^{\prime\prime})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\tfrac{\delta^{2}}{\ell^{2}}d_{2}(v_{1},p)^{2}}$. If $d_{2}(v_{1},p)\leq\alpha\ell$ then $d_{2}(p,p^{\prime\prime})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}\delta^{2}}\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}}$ since $\delta\leq 1$. By considering $p$ above as the location of $v_{i}$ ($1\leq i\leq\alpha\ell$) that maximizes its distance to $u_{i}$ (whose location corresponds to $p^{\prime\prime}$ above), we conclude that $d_{2}(v_{i},u_{i})\leq\sqrt{\lambda^{\prime 2}\gamma^{2}+\alpha^{2}}=\sqrt{(2\alpha-\alpha^{2})\gamma^{2}+\alpha^{2}}$, as desired. ∎